OpenStudy (anonymous):

Karen is financing $291,875 to purchase a house. She obtained a 15/5 balloon mortgage at 5.35%. What will her balloon payment be?

OpenStudy (anonymous):

So what I know is that the 15 means it'll be paid like it's going to be paid off in 15 years. However, they make those payments for 5 years since it's a 15/5. The remaining amount has tax added somewhere and then the remaining is the balloon. @rational

OpenStudy (anonymous):

Found something to work with for the monthly payment calculation

OpenStudy (rational):

you want to calculate monthly payment as if it is a 15 year loan ?

OpenStudy (anonymous):

Yes, but they pay it for 5 years

OpenStudy (anonymous):

Once we have the amount they paid over 5 years, we subtract that from the original amount.

OpenStudy (rational):

gotcha lets calculate monthly payment first then

OpenStudy (anonymous):

Yeah, once we have that the rest comes together pretty simply from what I can tell

OpenStudy (rational):

do we need to multiply it by 12*5 to get the amount paid ?

OpenStudy (anonymous):

2361.69 * 60 = 141,701.40

OpenStudy (anonymous):

Next we subtract that from the original mortgage

OpenStudy (anonymous):

$291,875 - 141,701.60 = $150,173.40

OpenStudy (rational):

that may not work, as "141,701.40" includes interest also.

OpenStudy (amistre64):

we dont make a full 5 years of payments, we make 1 less

OpenStudy (anonymous):

I believe its 5 years

OpenStudy (anonymous):

After you find the difference of original mortgage and payments made, my lesson says to apply one month's interest to the remaining and then you have the remaining balloon.

OpenStudy (amistre64):

payments are a geometric progression: P(k^n-1)/(k-1) defines the value of all payments made with respect to accruing interest the balance of the loan if Bk^n when the compounding loan is balanace by the payments, we have the loan paid off \[Bk^n=P\frac{k^n-1}{k-1}\] monthly payment P is just \[P=Bk^n\frac{k-1}{k^n-1}\]

OpenStudy (amistre64):

P = 291875k^(15*12) (k-1)/(k^(15*12)-1), k=1+.0535/12

OpenStudy (amistre64):

5*12 - 1 is the number of regular monthly payments, the remainder the ballon, is the final 60th payment

OpenStudy (amistre64):

recall that Bn=0 defines the ending balance, therefore \[B_n=Bk^n-P\frac{k^n-1}{k-1}\] \(B_{59}\) is our next to last payment, and then interest accrues for one more month \(B_{59}~k\) is our ballon payment

OpenStudy (amistre64):

any of this make sense?

OpenStudy (amistre64):

B0 = B , we take out a loan B1 = Bk - P , interest accrues and we make a payment B2 = B1 k - P , interest accrues and we make a payment = (Bk - P)k - P = Bk^2 - P - Pk B3 = B2 k - P , interest accrues and we make a payment = (Bk^2 - P - Pk)k - P = Bk^3 - P - Pk - Pk^2 ..... Bn = Bk^n - P(sum of a geometric sequence in k)

OpenStudy (amistre64):

when Bn = 0 0 = Bk^n - P(k^n-1)/(k-1) Bk^n = P(k^n-1)/(k-1) which is what i stated at the begining to define payment amounts

OpenStudy (anonymous):

Does this seem correct?

OpenStudy (anonymous):

im just wondering how your formula for payment coincides with the common PV payment formula \[P=\frac{PV*r}{1-(1+r)^{-n}}\]

OpenStudy (anonymous):

I see how you derived your formula ... and I know how this formula is derived But do they actually equal each other ? Or are they diff formulas?

OpenStudy (amistre64):

B_59 = 291875k^(59) - (2361.69)(k^59-1)/(k-1), k=1+.0535/12 B_59 = 220491.35 B_60 = B_59 k = ballon payment

OpenStudy (amistre64):

they are equal, i just hate the textbook version of it, it doesnt convey the mechanics of it and relies to heavily on trying to memorize a format

OpenStudy (anonymous):

oh ok cool

OpenStudy (amistre64):

PV is a tabled value isnt it ....

OpenStudy (amistre64):

i use my formula to find many many things that my accounting classes kept trying to give different formulas for ... why learn 12 different things, when 1 will do fine?

OpenStudy (anonymous):

LOL I guess so But the otehr derivation is very helpful when the payments vary every month or ya wanna find the net Present value

OpenStudy (amistre64):

depreciation was a hoot if there is no salvage value: use cost/life to define straight line depreciation if we have salvage value: use the formula (cost - salvage)/life why? why not define it as: (cost - salvage)/life to start with?

OpenStudy (amistre64):

i havent tried to generalize it to varying payments yet :) but yeah, that might be a pro for it

OpenStudy (amistre64):

on my the tests, while everyone was lookng up tables and trying to find the proper format of the formulas, i was just plugging in values and moving right along lol

OpenStudy (anonymous):

LOL i feel like id be the one scrambling for formulas K can we just clarify the balloon payment method since i am not sure i follow it So basically you divide the payments as if you were paying it over 15 years In actuality you only pay for 5 years Then the remainder of what is left over you pay in one lump sum Is it the remainder of payments? Or is the remainder of the loan?

OpenStudy (amistre64):

the monthly payments are defined for the usual fixed rate loan of say 15 years we pay off that loan like usual for 5 years, except for one difference the next to last payment is the last monthly payment, the 60 payment is the ballon, that amount of the after the last monthly payment, compounded for 1 more month

OpenStudy (amistre64):

59P + balloon = total cost of the loan

OpenStudy (amistre64):

how much do we save in interest? do we save in interest?

OpenStudy (amistre64):

12*15P - principal = interest (on a 15 year loan) 58P + balloon - principal = interest (on a 15/5 )

OpenStudy (anonymous):

Its 59 not 58 right? 59P + balloon - principal = interest (on a 15/5 )

OpenStudy (amistre64):

12*5 = 60 payments, 59 are monthly, 1 is balloony

OpenStudy (amistre64):

yeah, slip of the finger lol

OpenStudy (anonymous):

lol okkk

OpenStudy (amistre64):

12*15*2361.69 - 291875 = 133229 220491 + 221474 - 291875 = 150090 of course we would expect to pay more in interest for a higher risk loan. it favors the investor.

OpenStudy (anonymous):

RIght so less interest with balloon method

OpenStudy (amistre64):

nope, more interest

OpenStudy (anonymous):

waitttt

OpenStudy (anonymous):

the more compounded the less interest

OpenStudy (anonymous):

but why wld someone choose this balloon method then -.-

OpenStudy (anonymous):

It pays off the mortgage quickly

OpenStudy (amistre64):

15 years has a balance after 5 years ... we havent reduced the balance enough to compensate

OpenStudy (amistre64):

why does anyone take a higher interest loan? desperation of course

OpenStudy (anonymous):

Got stuck on another question dealing with "weighted mean". You guy's willing to help?

OpenStudy (amistre64):

link to the post?

OpenStudy (anonymous):

K thanks guys ... gonna chew over this

OpenStudy (anonymous):

@amistre64 I thought about your statement earlier that "more interest is accrued when using the balloon method" and I kinda disagree. The longer the loan is unpaid the more interest it accrues. Every compounding period we have to pay interest for the amount that is still outstanding. Therefore with the Balloon method you pay less interest. 12*15*2361.69 - 291875 = 133229 --> 15 Year Loan 139339.71 + 221474 - 291875 = 68938.71 ---> Balloon Method

OpenStudy (amistre64):

pfft, 220491 was the balance that accrued one more interest for balloon. i used the wrong value .... i was thinking along the lines that the payments mostly cover interest in the beginning and not so much the principal. I wasnt even thinking that i had an error :)

OpenStudy (amistre64):

i do like being proved wrong ;) keeps me learning lol

OpenStudy (anonymous):

hahaha You kept me pondering over this alll evening. Was killing my brain